# Can World Simulators Reason? Gen-ViRe: A Generative Visual Reasoning Benchmark

### v2

@@ 1 Introduction
The potential of this paradigm is evidenced by rapid progress in video generation technology. Large-scale models—from commercial systems like Sora [2] and Kling [28] to open models like CogVideo [19] and HunyuanVideo [27]—have demonstrated impressive emergent capabilities in understanding physical interactions, temporal causality, and spatial relationships. Qualitative demonstrations from pioneering works are compelling: models generating coherent multi-step processes, respecting object permanence, and exhibiting intuitive physics. Yet despite this progress, a fundamental question remains unanswered: how well do these models actually reason? Qualitative showcases, while impressive, provide no systematic measure of reasoning depth. More critically, existing video generation benchmarks, which focus primarily on fidelity, alignment, or object-level consistency, do not assess CoF reasoning and therefore cannot measure a model’s core cognitive abilities in multi-step planning and algorithmic logic, or abstract pattern extrapolation [20, 33, 32, 11, 18, 39, 42, 47, 40, 30, 6, 29, 10, 50]. Without rigorous quantitative evaluation, we cannot distinguish genuine reasoning from sophisticated pattern matching, nor can we diagnose where models fail—perception, physics understanding, or planning—to guide systematic improvement.
@@
To address this critical evaluation gap, we propose Gen-ViRe (Generative VisualReasoning Benchmark), the first comprehensive framework purpose-built to systematically assess CoF reasoning by video generation models. Gen-ViRe is grounded in dual foundations: cognitive science principles defining core pillars of human reasoning [37, 26, 14, 25, 36, 38, 35, 4] , and application-driven requirements from emerging domains like embodied AI and autonomous systems [12, 49]. We decompose Generative VisualReasoning into six fundamental dimensions— Perceptual, Analogical, Algorithmic & Logical, Spatio-Temporal & Dynamic, Procedural & Planning, and Abstract Reasoning—spanning the complete spectrum from foundational perception to high-order planning [31, 48, 1]. Our data curation employs a multi-source strategy as shown in Figure 4: curating from web resources and academic papers, integrating existing datasets (e.g., ARC-AGI for abstract reasoning, KiVA for analogical tasks), and generating novel tasks using T2I models where suitable data is absent. We implement rigorous prompt design through minimal prompting principles—deliberately providing only high-level goals to assess autonomous reasoning rather than instruction-following—validated through iterative peer review. Our evaluation methodology features a hybrid VLM-assisted approach: Image VLMs judge final outputs for static reasoning tasks, while Video VLMs assess the complete generation process for dynamic tasks, with each judge provisioned with detailed, sub-category-specific criteria refined through multi-round human validation. Through extensive experiments evaluating 7 state-of-the-art video generation models across 72 distinct reasoning prompts (with each model generating 5 instances per prompt, totaling 360 videos per model and over 2,500 overall), Gen-ViRe delivers the first systematic quantification of CoF reasoning capabilities, revealing critical gaps between qualitative potential and quantitative performance, and establishing scientific baselines to guide the development of genuine world simulators.
@@ 3.2 Taxonomy of Generative Visual Reasoning
Our proposed Gen-ViRe taxonomy is rooted in two complementary perspectives: (i) the theoretical foundations of cognitive science and (ii) the practical demands of key emerging GVR application domains. More specifically: (i) Cognitive Science Foundations: A robust GVR benchmark should be able to evaluate the core pillars of human cognitive ability [4, 11, 19, 20, 27–30]. This provides the theoretical basis for our categories of Perceptual Reasoning, Analogical Reasoning, and Abstract Reasoning. (ii) Application-Driven Requirements: Concurrently, as AI expands from the symbolic to the physical realm, emerging applications (e.g., Embodied AI, Autonomous Driving) demand dynamic, multi-step reasoning [9, 36]. These capabilities are essential for the next frontier of AI but cannot be evaluated by existing benchmarks. These practical requirements provide the basis for our categories of Planning Reasoning, Spatial/Temporal Reasoning, and Algorithmic & Logical Reasoning.

### v3

@@ 1 Introduction
The potential of this paradigm is evidenced by rapid progress in video generation technology. Large-scale models—from commercial systems like Sora [2] and Kling [28] to open models like CogVideo [20] and HunyuanVideo [26]—have demonstrated impressive emergent capabilities in understanding physical interactions, temporal causality, and spatial relationships. Qualitative demonstrations from pioneering works are compelling: models generating coherent multi-step processes, respecting object permanence, and exhibiting intuitive physics. Yet despite this progress, a fundamental question remains unanswered: how well do these models actually reason? Qualitative showcases, while impressive, provide no systematic measure of reasoning depth. More critically, existing video generation benchmarks, which focus primarily on fidelity, alignment, or object-level consistency, do not assess CoF reasoning and therefore cannot measure a model’s core cognitive abilities in multi-step planning and algorithmic logic, or abstract pattern extrapolation [21, 30, 31, 11, 19, 34, 27, 10]. Without rigorous quantitative evaluation, we cannot distinguish genuine reasoning from sophisticated pattern matching, nor can we diagnose where models fail—perception, physics understanding, or planning—to guide systematic improvement.
@@
To address this critical evaluation gap, we propose Gen-ViRe (Generative VisualReasoning Benchmark), the first comprehensive framework purpose-built to systematically assess CoF reasoning by video generation models. Gen-ViRe is grounded in dual foundations: cognitive science principles defining core pillars of human reasoning [25, 15, 24, 4, 33, 5] , and application-driven requirements from emerging domains like embodied AI and autonomous systems [12, 40]. We decompose Generative VisualReasoning into six fundamental dimensions— Perceptual, Analogical, Algorithmic & Logical, Spatio-Temporal & Dynamic, Procedural & Planning, and Abstract Reasoning—spanning the complete spectrum from foundational perception to high-order planning [1]. Our data curation employs a multi-source strategy as shown in Figure 4: curating from web resources and academic papers, integrating existing datasets (e.g., ARC-AGI for abstract reasoning, KiVA for analogical tasks), and generating novel tasks using T2I models where suitable data is absent. We implement rigorous prompt design through minimal prompting principles—deliberately providing only high-level goals to assess autonomous reasoning rather than instruction-following—validated through iterative peer review. Our evaluation methodology features a hybrid VLM-assisted approach: Image VLMs judge final outputs for static reasoning tasks, while Video VLMs assess the complete generation process for dynamic tasks, with each judge provisioned with detailed, sub-category-specific criteria refined through multi-round human validation. Through extensive experiments evaluating 7 state-of-the-art video generation models across 72 distinct reasoning prompts (with each model generating 5 instances per prompt, totaling 360 videos per model and over 2,500 overall), Gen-ViRe delivers the first systematic quantification of CoF reasoning capabilities, revealing critical gaps between qualitative potential and quantitative performance, and establishing scientific baselines to guide the development of genuine world simulators.
@@ 3.2 Taxonomy of Generative Visual Reasoning
Our proposed Gen-ViRe taxonomy is rooted in two complementary perspectives: (i) the theoretical foundations of cognitive science and (ii) the practical demands of key emerging GVR application domains. More specifically: (i) Cognitive Science Foundations: A robust GVR benchmark should be able to evaluate the core pillars of human cognitive ability [7, 1, 5, 15, 24, 25, 33],. This provides the theoretical basis for our categories of Perceptual Reasoning, Analogical Reasoning, and Abstract Reasoning. (ii) Application-Driven Requirements: Concurrently, as AI expands from the symbolic to the physical realm, emerging applications (e.g., Embodied AI, Autonomous Driving) demand dynamic, multi-step reasoning [17]. These capabilities are essential for the next frontier of AI but cannot be evaluated by existing benchmarks. These practical requirements provide the basis for our categories of Planning Reasoning, Spatial/Temporal Reasoning, and Algorithmic & Logical Reasoning.